OpenStudy (anonymous):

82 1/2% of what number is 33?

OpenStudy (anonymous):

\[82 \frac{ 1 }{ 2}\]

OpenStudy (slaaibak):

you can solve this by writing the equation: x * 82.5/100 = 33 so x = 33 * 100 / 82.5 = 40

OpenStudy (anonymous):

thank you @slaaibak